# CAOTE: KV Cache Eviction for LLMs via Attention Output Error-Based Token Selection

### v4

@@ Abstract
While long context support of large language models has extended their abilities, it also incurs challenges in memory and compute which becomes crucial bottlenecks in resource-restricted devices. Token eviction, a widely adopted post-training methodology designed to alleviate the bottlenecks by evicting less important tokens from the cache, typically uses attention scores as proxy metrics for token importance. However, one major limitation of attention score as a token-wise importance metrics is that it lacks the information about contribution of tokens to the attention output. In this paper, we propose a simple eviction criterion based on the contribution of cached tokens to attention outputs. Our method, CAOTE, optimizes for error due to token eviction, by seamlessly integrating attention scores and value vectors. This is the first method to use information from the value vector on top of attention-based eviction scores. Additionally, CAOTE can act as a meta-heuristic method with flexible usage with any token eviction method. We show that CAOTE, when combined with state-of-the-art attention score-based methods, always improves accuracies on the downstream task for LLAMA3 and QWEN2.5 model families, indicating the importance of leveraging information from values during token eviction process.
@@ 1 Introduction
Large Language Models (LLMs) represent a large step forward in natural language processing by demonstrating remarkable proficiency in tasks such as text generation (Coenen et al., 2021), machine translation (Xiao et al., 2023), and question-answering (Robinson et al., 2022). Many of these tasks require handling long prompt inputs efficiently, particularly in retrieval-augmented generation (RAG), long-form document understanding (Liao et al., 2024), summarization (Zhang et al., 2024a), and multi-turn dialogue systems (Thoppilan et al., 2022)—collectively referred to as long-context LLMs. A key challenge in long-context applications is the increased latency during inference and generation, primarily due to the quadratic computational complexity of self-attention and the growing memory demands of handling long sequences.
To alleviate computational overhead, Key-Value (KV) caching is a widely adopted technique that enables faster inference in LLMs. It does so by storing the key-value states of previously processed tokens and reusing them when generating new tokens, reducing redundant computations (Pope et al., 2023). However, while KV caching significantly enhances efficiency, it comes at the cost of substantial memory consumption as mentioned in (Zhang et al., 2024b) , especially in long-context setups. In such cases, the memory footprint of the KV cache often exceeds the memory usage of the model itself, making it a major bottleneck, especially, for memory-constraint devices. In this paper we focus on improving KV cache eviction methods by keeping relevant KV to utilize the memory budget in the best possible way.
Several methods have been proposed to optimize KV cache memory usage, including sparse attention mechanisms that selectively attend to a subset of tokens rather than all previous tokens reducing memory and computational complexity (Zhang et al., 2023; Xiao et al., 2024), efficient attention architectures such as linearized attention (Choromanski et al., 2020), and memory-efficient transformers (Shazeer, 2019) which approximate self-attention to minimize memory consumption. These methods require from-scratch training or model finetuning, however, in this work, we focus on post-training token eviction method which dynamically discard less important tokens from the KV cache to control memory growth (Zhang et al., 2024b). Token eviction methods offer a distinct advantage in memory management by explicitly reducing the KV cache size while preserving critical information. These methods leverage the sparsity of attention — a phenomenon where a small subset of cached keys contribute disproportionately in the attention mechanism — to selectively retain only the most important tokens.
Typically, token importance is determined based on attention scores or, equivalently, attention weights, which measure the alignment between query and key tokens (Zhang et al., 2024b; Oren et al., 2024). However, since the attention output is a linear combination of attention weights and value vectors, evicting tokens simply based on attention scores may lead to suboptimal decisions, as it does not consider the contribution of value vectors from the tokens to be evicted.
In this paper, we propose CAOTE, a post-training cache eviction method that seamlessly integrates eviction (attention) scores with value vectors by optimizing the eviction error, unlike recent methods, as shown in Table 1. CAOTE offers flexibility by being applicable to any post-training eviction method and also includes an efficient counterpart. When combined with recent cache eviction methods (Zhang et al., 2024b; Oren et al., 2024; Li et al., 2024), we observe performance boost for all recent token eviction methods on a variety of downstream tasks: 16 tasks from LongBench, Needle-in-Haystack and Perplexity measurement tasks.
@@ 4.2 LongBench
We present the accuracy of Llama 3.1-8B-Instruct, Llama 3.2-3B-Instruct and, Qwen 2.5-3B-Instruct, Qwen 2.5-7B-Instruct using baseline eviction methods with budget of 22k, 44k, both with and without CAOTE in Table 2 and Table 3. We observe that the best average performance is given by SnapKV-FastCAOTE for the Llama3 models, while for Qwen 2.5 models SnapKV-CAOTE performs the best. H2O shows $30\%$ improvement with CAOTE, while TOVA, SnapKV also show overall improvements, making their average accuracy closer to dense accuracy. Additionally, for some QA tasks (Qasper, MF-en, Musique, 2WikiMQA), H2O-(Fast)CAOTE performs best. Additional results for the $6$ k, $8$ k budget are shown in Table 9, Table 10 for Llama3 and Qwen 2.5 respectively, in Section E.1, which follow a trend similar to the 22k, 44k budgets.

### v5

# CAOTE: KV Cache Selection for LLMs via Attention Output Error-Based Token Eviction

@@ Abstract
While long context support of large language models has extended their abilities, it also incurs challenges in memory and compute which becomes crucial bottlenecks in resource-restricted devices. Token eviction, a widely adopted post-training methodology designed to alleviate the bottlenecks by evicting less important tokens from the cache, typically uses attention scores as proxy metrics for token importance. However, one major limitation of attention score as a token-wise importance metrics is that it lacks the information about contribution of tokens to the attention output. In this paper, we propose a simple eviction criterion based on the contribution of cached tokens to attention outputs. Our method, CAOTE, optimizes for error due to token eviction, by seamlessly integrating attention scores and value vectors. This is the first method to use information from the value tokens on top of attention-based eviction scores in closed-form. Additionally, CAOTE can act as a meta-heuristic method with flexible usage with any token eviction method. We show that CAOTE, when combined with state-of-the-art attention score-based methods, always improves accuracies on the downstream task for LLAMA3 and QWEN2.5 model families, indicating the importance of leveraging information from values during token eviction process.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive capabilities across a wide range of natural language tasks, including text generation (Coenen et al., 2021), machine translation (Xiao et al., 2023), and question answering (Robinson et al., 2022). Many of these applications—such as retrieval-augmented generation (RAG), long-form document understanding (Liao et al., 2024), summarization (Zhang et al., 2024a), and multi-turn dialogue systems (Thoppilan et al., 2022)—require processing long input sequences, giving rise to the class of long-context LLMs.
A central challenge in long-context inference is the increased computational and memory overhead, particularly during the prefill and generation phases. This is primarily due to the quadratic complexity of self-attention and the growing memory footprint from storing activations over extended sequences. To mitigate quadratic attention computation, Key-Value (KV) caching has become a standard optimization, enabling faster inference by reusing previously computed key-value states (Pope et al., 2023). However, in long-context settings, the memory consumed by the KV cache can surpass that of the model itself (Zhang et al., 2024b), posing a significant bottleneck—especially for memory-constrained environments.
Recent efforts to manage KV cache memory can be broadly categorized into training-based approaches (Zhang et al., 2023; Xiao et al., 2024) and post-training methods. This work focuses on the latter, which operate under fixed memory budgets and dynamically evict tokens from the cache. Early post-training methods such as H2O (Zhang et al., 2024b) retain tokens with the highest cumulative attention scores, while subsequent approaches refine this strategy using variants of attention-based metrics (Oren et al., 2024; Li et al., 2024; Tang et al., 2024; Qin et al., 2025). These methods exploit the sparsity of attention—where a small subset of tokens disproportionately influence the output—to preserve only the most salient cached entries.
A key limitation of existing eviction methods is their reliance solely on attention scores, which reflect the alignment between query and key tokens. However, the output of the self-attention layer—the attention output—is a weighted combination of these scores and the corresponding value tokens. Since this output directly influences the hidden state and ultimately the model’s predictions, ignoring the contribution of value tokens can lead to suboptimal eviction decisions.
In this paper, we propose CAOTE (Cache Selection via Attention Output error-based Token Eviction), a post-training eviction method that directly computes the impact of each token on the attention output in closed form. Unlike prior approaches, CAOTE integrates both key and value contributions to estimate the exact eviction error during generation as shown in Table 1. It is model-agnostic and can be applied as a meta-eviction strategy atop existing score-based methods. We also propose an efficient variant of CAOTE suitable for deployment in constrained environments.
Empirically, CAOTE consistently improves performance across a diverse set of benchmarks, including 16 tasks from LongBench, Needle-in-Haystack retrieval, and perplexity evaluations. When combined with recent eviction strategies (Zhang et al., 2024b; Oren et al., 2024; Li et al., 2024), CAOTE yields substantial gains in accuracy and efficiency.
@@ 4.2 LongBench
We present the accuracy of Llama 3.1-8B-Instruct, Llama 3.2-3B-Instruct and, Qwen 2.5-3B-Instruct, Qwen 2.5-7B-Instruct using baseline eviction methods with budget of 22k, 44k, both with and without CAOTE in Table 2 and Table 3. We observe that the best average performance is given by SnapKV-FastCAOTE for the Llama3 models, while for Qwen 2.5 models SnapKV-CAOTE performs the best. H2O shows $30\%$ improvement with CAOTE, while TOVA, SnapKV also show overall improvements, making their average accuracy closer to dense accuracy. Moreover, we obserive that for some QA tasks (Qasper, MF-en, Musique, 2WikiMQA), H2O-(Fast)CAOTE performs best.
Additional results for the $6$ k, $8$ k budget are shown in Table 9, Table 10 for Llama3 and Qwen 2.5 respectively, in Section E.1, which follow a trend similar to the 22k, 44k budgets.